OpenStudy (anonymous):

Having trouble with how to do this problem Find the point on the line -4 x + 5 y - 4 =0 which is closest to the point ( -1, -3 )

OpenStudy (dumbcow):

a point is always closest to a line when at a 90 degree angle find a line perpendicular to given line that goes through point (-1,-3) then solve for the point where 2 lines intersect

OpenStudy (dumbcow):

first get line in slope-intercept form \[y = \frac{4}{5} x + \frac{4}{5}\] slope = 4/5 perpendicular slope = -5/4 find equation of line through point (-1,-3) using point slope form \[y-(-3) = -\frac{5}{4}(x -(-1))\] \[y = -\frac{5}{4}x -\frac{17}{4}\] set lines equal \[\frac{4}{5}x + \frac{4}{5} = -\frac{5}{4}x - \frac{17}{4}\] solve for x ...

OpenStudy (anonymous):

awesome thanks!!!!

OpenStudy (dumbcow):

yw